OpenStudy (anonymous):

the mass % and mole fraction of an unknown solute are 21% and .022 respectively. What is the molar mass of the solute. I've done this problem before but now I'm stuck on this....

OpenStudy (anonymous):

Preferably a hint would be better than completely working it out

OpenStudy (anonymous):

I don't think you can do it. You've got two equations: \[\frac{m}{m + m_s} = 0.21\] and \[\frac{\frac{m}{A}}{\frac{m}{A} + \frac{m_s}{A_s}} = 0.022\] But four unknowns: m (mass of solute), m_s (mass of solvent), A (molar mass of solute) and A_s (molar mass of solvent).

OpenStudy (anonymous):

yeah I'm starting to think that as well but I could've sworn I did it previously. I've been analyzing it and it doesn't seem do-able

OpenStudy (anonymous):

Well, try this. Pick out any m and m_s you want, then use the equations to calculate A and A_s. If you can do that and make it all work, then you've proved that you cannot find a unique solution -- because you can make the equations all work after choosing completely arbitrary values for two of the unknowns.

OpenStudy (anonymous):

Yeah I don't think its possible lol

OpenStudy (anonymous):

nvm its possible so .21= (.022x)/((.978molesH2O/18)+.022x) stupidest thing I've ever seen in my life....

OpenStudy (anonymous):

Ah. You didn't mention you knew the solvent was water. Key information.